Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability of rain

Expert replies
by wilderness » Fri Jul 25, 2008 9:56 am
A question with answer is given below.

How can we solve if the probability of rain on a given day is not 50%. Let say the queston was

"If the probability of rain on a certain day is 20%, what is the probability that it will rain in exactly 3 out of a 5-day period ? "

Will we only divide our current results by 2.5 since 50/20 is 2.5 or is there some other method.
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by Ian Stewart » Fri Jul 25, 2008 10:31 am
That makes the problem a bit more difficult. Say, first, you wanted to know the probability of rain on the first three days, then no rain on the last two days, and the probability of rain is 0.2. Then the probability would be:

(0.2)^3 * (0.8)^2

Of course, the probability will be the same for any other particular sequence of three rainy days and two non-rainy days. As in the problem above, there are 5C3 = 10 different sequences, so we need to multiply the above by 10.

10* (0.2)^3 * (0.8)^2 should be the answer, if you change the probability of rain to 0.2. Mind you, I'd be a bit surprised to see that on a GMAT- it's not completely out of the question, but it's unlikely.
For online GMAT math tutoring, or to buy my higher-level Quant books and problem sets, contact me at ianstewartgmat at gmail.com

ianstewartgmat.com
Join the discussion

I am still confused

by saabidi » Wed Nov 26, 2008 1:05 pm
Hi Guys,

Can you please explain how # of desired outcome is being calculated.
I am having hard time understanding how the answer is 10?

Thanks
Join the discussion

I am still confused

by saabidi » Wed Nov 26, 2008 1:07 pm
Hi Guys,

Can you please explain how # of desired outcome is being calculated.
I am having hard time understanding how the answer is 10?

Thanks
Join the discussion

by muzali » Wed Nov 26, 2008 2:58 pm
Probability of rain on any day=50% = 1/2
Probability of rain on 3 days out of 5 i.e. it rains on 3 days and does not rain on 2 days= [(1/2)^3]*[(1/2)^2] = (1/2)^5=1/32

Number of ways you can arrange these three days on which it will rain out of a total of five days = 5C3 = 5*4/2 = 10

Therefore, the required probability is = (1/32)*10 = 10/32=5/16
Join the discussion

Re: Probability of rain

by muzali » Sat Nov 29, 2008 5:03 pm
wilderness wrote:A question with answer is given below.

How can we solve if the probability of rain on a given day is not 50%. Let say the queston was

"If the probability of rain on a certain day is 20%, what is the probability that it will rain in exactly 3 out of a 5-day period ? "

Will we only divide our current results by 2.5 since 50/20 is 2.5 or is there some other method.
Sorry, had not seen the 20% question...
the appraoch is the same as 50%

Probability of rain on any day=20% = 1/5
Probability of rain on 3 days out of 5 i.e. it rains on 3 days and does not rain on 2 days= [(1/5)^3]*[(1/5)^2] = (1/5)^5

Number of ways you can arrange these three days on which it will rain out of a total of five days = 5C3 = 5*4/2 = 10

Therefore, the required probability is = ((1/5)^5)*10 = 2/5^4 =2/625
Join the discussion

Re: Probability of rain

by jimmiejaz » Mon Dec 01, 2008 6:22 am
muzali wrote:
wilderness wrote:A question with answer is given below.

How can we solve if the probability of rain on a given day is not 50%. Let say the queston was

"If the probability of rain on a certain day is 20%, what is the probability that it will rain in exactly 3 out of a 5-day period ? "

Will we only divide our current results by 2.5 since 50/20 is 2.5 or is there some other method.
Sorry, had not seen the 20% question...
the appraoch is the same as 50%

Probability of rain on any day=20% = 1/5
Probability of rain on 3 days out of 5 i.e. it rains on 3 days and does not rain on 2 days= [(1/5)^3]*[(1/5)^2] = (1/5)^5

Number of ways you can arrange these three days on which it will rain out of a total of five days = 5C3 = 5*4/2 = 10

Therefore, the required probability is = ((1/5)^5)*10 = 2/5^4 =2/625
Muzali,
Just to correct you....
The probability that it does not rain on any day is 1-1/5 = 4/5
so, as Ian explained above probability will be
[(1/5)^3]*[(4/5)^2]*10
hope it helps...
What if i have not yet beat the beast, I know i will beat it!!!!!!!!
Join the discussion

by tohellandback » Tue Jun 30, 2009 11:07 pm
muzali wrote:Probability of rain on any day=50% = 1/2
Probability of rain on 3 days out of 5 i.e. it rains on 3 days and does not rain on 2 days= [(1/2)^3]*[(1/2)^2] = (1/2)^5=1/32

Number of ways you can arrange these three days on which it will rain out of a total of five days = 5C3 = 5*4/2 = 10

Therefore, the required probability is = (1/32)*10 = 10/32=5/16
I have seen this answer in different threads but still don't understand why do you calculate the arrangement and then multiply.
could you please explain?
The powers of two are bloody impolite!!
Join the discussion

by Robinmrtha » Tue Jun 30, 2009 11:30 pm
the probability of raining is 1/2
so the probability that it will rain exactly three days in five days will be
rains in 1st day x rains in 2nd day x rains in 3rd day x doesnt rain in 4th day x doesnt rain in 5th day + ..........(all possible arrangements of three)
i.e. total ways 3 days can be selected from 5
or 5C3=10
now since the probability is half...
the probability for raining and not raining both is same i.e. 1/2
so
so for any combination of days the probability will be same
i.e 1/2x1/2x1/2x1/2x1/2=1/32
And probability will repeat for all possible combination. i.e. 10
so multiply by 10
i.e. 10x 1/32 =5/16
hope this helps
Join the discussion

by goalevan » Mon Mar 01, 2010 6:27 pm
The binomial distribution formula comes in handy for this problem. Basically what most of the previous posts have contained:

p Probability of success (rain): 1/5
n Number of trials: 5
k Number of successes: 3

Binomial distribution formula: nCk * p^k * (1-p)^(n-k)

In other words, the possible combinations of k successes in n trials (different ways to order those k items) multiplied by the probability of one of those strings occurring.

nCk = 10
p^k = (1/5)^3 = 1/125
(1-p)^(n-k) = (4/5)^2 = 16/25
10 * 1/125 * 16/25 = 32/625 = 5.12%
Join the discussion